# Machine-learning inference of stellar properties using integrated photometric and spectroscopic data

### v1

@@ 4.1 Spectra preprocessing
LAMOST is a low-resolution ($\frac{\lambda}{\Delta\lambda}\sim 1800$) spectrometer with two arms: a blue arm that covers a wavelength range of $3700-5900$Å, and a red arm that covers a wavelength range of $5700-9000$Å. The LAMOST survey consists of millions of spectra of stars, galaxies, and quasars. We used LAMOST DR8 and downloaded all the spectra of $A,F,G,K,M$ stars with $3000\leq T_{\mathrm{eff}}\leq 7500$K. We removed samples without measured $T_{\mathrm{eff}},\log g$ or $[\mathrm{Fe}/\mathrm{H}]$ (from the LASP (Wu et al., 2014) pipeline), and samples with nonsensical negative SNR. This resulted in about $6.5$M samples. As mentioned in Section 3.1, the hybrid approach does not require having supervised labels. Here, we decided to use only samples with labels, given the available large size of the dataset. The preprocessing of spectra is similar to what was suggested in StarGRUNet (Li & Lin, 2023): we first translate the wavelength to the rest-frame using radial velocity measurements from the LASP pipeline (Wu et al., 2014). Next, we divide the spectrum into blue and red regions and resample each region using linear interpolation. We then apply a median filter to each region, followed by a continuum normalization using a fifth-order polynomial, and a final step where we remove points higher or lower than ±$\pm 3\sigma$ and normalize the flux to have zero mean and unit standard deviation. For a more detailed explanation of the different pre-processing steps, please refer to Li & Lin (2023). Figure 3 shows an example of the pre-processing stages of LAMOST spectra.
@@ 4.2 Light curve preprocessing
Kepler was a space mission designed to provide high cadence light curves for stars and for the search of exoplanets through transits. As such, Kepler measured the light curves of around $200000$ main sequence and giant stars for almost $4$ years with a cadence of approx. $30$ minutes. We used Kepler samples with stellar properties from Berger et al. (2020), resulting in $183,435$ samples. The preprocessing of the light curves follows two different ways: one using the absolute magnitude and the other using the mean and standard deviation. The reason for the two normalizations is that different properties require different light curve information. The activity and luminosity, for example, are related to the absolute amplitude of the light curve (normalized by the distance). On the other hand, period is more easily detected using standard normalization, like zero mean and unit standard deviation. The absolute magnitude normalization was done by first dividing the light curve by $2^{-k}$ where $k$ is the absolute magnitude, calculated using the Kepler magnitude, KMAG, and the distance from Berger et al. (2020). In addition to the raw light curve, as mentioned earlier, we also calculated the Fast Fourier Transform (FFT) and the auto correlation function (ACF), which help to determine the period. Figure 4 shows an example of the pre-processing stages of the Kepler light curve.
@@ 5 Implementation details
In big models, like DESA, hyperparameter tuning can be a very challenging task. This becomes even harder when the model consists of two steps - pre-training and alignment. We therefore chose to use a simple heuristic when defining the hyperparameters of our model. As the number of spectra samples is much larger than the number of light curve samples ($6.5$M vs $200$K), we designed the individual encoders such that the spectra encoder has more parameters than the light curve encoder. This is motivated by neural scaling laws, a phenomenological relationship between dataset size, model parameters, and performance that was originally found in the vision and language domains (Kaplan et al., 2020), but recent works showed it also applies to astronomical data (Walmsley et al., 2024; Pan et al., 2024a). Therefore, the dimension of the final spectra features was chosen to be $2048$, and that of the light curve was chosen to be $256$. The number of parameters in the spectra and light curve encoders is about $500$M and $11$M, respectively. During hybrid training, $\lambda$ was chosen arbitrarily to be $0.5$. The embedding dimension in dualformer was set to $256$, and the number of parameters in this module is also about $11$M. The light curve encoder was trained using a learning rate decay scheduler with the cosine annealing method. The initial learning rate is $2\cdot 10^{-5}$ decreasing to $2\cdot 10^{-6}$. All other modules were trained with a constant learning rate of $2\cdot 10^{-5}$. We trained all modules with AdamW optimizer (Loshchilov & Hutter, 2017). Lastly, we estimated the energy used to train the entire model using CodeCarbon [^1] https://codecarbon.io/ package. It is estimated to be $334$ kWh for the entire model, out of which $204$ kWh are for the pretraining stage. All the code used for training and experiments is publicly available on https://github.com/IlayMalinyak/DESA.
@@ 6.1 Pretraining Results
First, we present the results of the hybrid pre-training of individual modalities. As mentioned before, the spectra encoder was trained using $T_{\mathrm{eff}},\log g$ and $[\mathrm{Fe}/\mathrm{H}]$ labels from the LASP pipeline. However, we test the results using labels from the high-resolution APOGEE survey (Abdurro’uf et al., 2022), at the same temperature range. This is a common practice and was done in previous works such as StarGRUNet (Li & Lin, 2023). The results on the test set of both LASP labels and APOGEE labels are shown in Figure 5. For each label, we also plot the $80\%$ and $50\%$ confidence intervals. It can be seen that the mean average error (MAE) on LASP is a factor of $\sim 2$ lower compared to APOGEE for all labels. This is reasonable given the fact that the APOGEE labels come from high-resolution spectra, while the spectra that are given to the model are low resolution. Another interesting difference between the labels is related to the prediction intervals. As mentioned in section 3.1, the prediction intervals were calibrated on a held-out validation set, different from the test set. This validation set uses the LASP labels. We see that for the LASP test set, this calibration results in the desired coverage ($80\%$ of the points are inside the $80\%$ interval, for example), while for the APOGEE labels, the coverage is lower. This is another evidence that there is a significant difference between the labels. We also compare our results with StarGRUNet. StarGRUNet (Li & Lin, 2023) used a GRU-based model to predict stellar parameters on a wide range of SNR values (SNR $>5$). While there are works with better performance for a specific range of SNR (Li et al. (2022), for example, who used $10<$ SNR $<20$), this is the state-of-the-art for a model that uses LAMOST low-resolution spectra and a wide range of SNR values.
Table 1 shows a comparison between StarGRUNet and DESA. We note that the results are very close. Specifically, our model performs slightly better than StarGRUNet for $T_{\mathrm{eff}}$ and $[\mathrm{Fe}/\mathrm{H}]$, and slightly worse for $\log g$. However, it is worth mentioning that we used SNR $>0$ and StarGRUNet used SNR $>5$. This seemed to be a result of the fact that we incorporated the SNR information into the training. In Appendix A we show MAE vs different SNRs bins for $T_{\mathrm{eff}},\log g,$ and $[\mathrm{Fe}/\mathrm{H}]$. We see that the sensitivity of our results to SNR is indeed better than StarGRUNet (see Figure 11 in their paper). However, there is still non-negligible sensitivity for both LASP and APOGEE labels. While it is probably impossible to remove all the dependency on the noise, this might be further investigated and improved in future papers.
@@
Model $T_{\mathrm{eff}}$ MAE (K) $\log g$ MAE (dex) $[\mathrm{Fe}/\mathrm{H}]$ MAE (dex)
@@ 6.3 Binary Detection
We fine-tuned our model, with an added classification head, for a binary classification of binarity - all the binaries were assigned as class $1$, regardless of the detection method, and all the singles were assigned as class 00. As baselines, we trained the same models that were used in 6.2 as well as unimodal models. The unimodal models are simply the pretrained encoders with an added prediction head. Using the pretrained encoders is important as a sanity check - we want to see if our model is able to extract meaningful information from the combination of modalities and, specifically, better than each modality alone. In all models, we made sure that the number of trainable parameters is not less than that in our DESA model ($11$M). Figure 12 compares the results using a precision-recall curve (right panel) and a True Negative Rate (TNR) vs. False Negative Rate (FNR) curve (left panel). Notice that a TNR-FNR curve can be seen as a flipped ROC curve, and therefore, the area under the curve (AUC) metric has the same meaning (with the area calculated in the opposite direction). Note that we opted for the TNR-FNR curve instead of the standard ROC curve for visual clarity. It can be seen that our model performs significantly better compared to all other models, with an AUC of $0.99$ and average precision (AP) of $1.00$. Figure 13 shows the confusion matrix of our model. It is interesting to compare it with the confusion matrices of all other models. Such a comparison revealed that, except for VicReg, all other models collapsed to predict the majority class (binary class). With the exception of the proposed DESA, VicReg was the only model able to predict some single stars correctly, but it did so with much lower precision for that class, in comparison to our model. The precision of VicReg on the single-star class is $48\%$, while our model achieved a precision of $88\%$ on that class. Confusion matrices of all other models are reported in Appendix 18.
@@ 6.4 Stellar age predicition
Model zero-shot accuracy Linear Regression $R^{2}$ $BP-RP$ Accuracy $Gmag$ Accuracy
@@
7 Conclusions
We presented DESA, a new multi-modality model for stellar astrophysics. DESA backbone consists of pre-trained unimodality encoders that were trained in a hybrid approach and shows state-of-the-art performance. The alignment module, DualFormer, is motivated by the observation that astrophysical data is unique and different compared to common multi-modality domains such as vision and NLP. We show that DESA learns a latent space that distills important physical information and can be easily transformed into meaningful diagrams. These diagrams can be used for tasks such as population analysis and outlier detection. The effectiveness of DESA is further demonstrated in various ways, including zero-shot, few-shot, and fine-tuning experiments to predict challenging and important labels such as binarity and age. DESA consistently outperforms all baselines on all experiments with (sometimes very) large margins, proving its superiority in the astronomical domain and marking a new era in data-driven astronomy where multiple measurements are efficiently combined to extract new insights. Each fine-tune task is of great physical importance and deserves an in-depth investigation and separate papers, which we plan for future work. DESA faces two primary limitations: interpretability and single-resolution modeling. Interpretability remains a challenge across machine learning models, lacking robust methods for results interpretation. Regarding resolution, DESA, while multi-modal, was trained on single surveys per modality (Kepler and LAMOST), without explicit consideration for survey-specific resolutions. Addressing interpretability and multi-resolution capabilities stands as our main goal for future research. Nonetheless, DESA represents a potential groundbreaking advancement in multimodal stellar astronomy.
@@ Appendix A Supplementary graphs
Figure 15: UMAP of all models. All points are in shaded gray, and 10 stars with pairs of spectra are colored on top. For each pair, all samples are colored with the same color and connected by a line. We can see the visual difference between DESA (upper figure) and all other models – in DESA, pairs arrange very close to each other, implying that they have very similar embeddings. In other models, there are samples of the same star that are very distant, which implies very different embeddings.
@@
Figure 17: UMAP of DESA final embeddings. All points are in gray, and points with RUWE $>1.4$ are colored according to their $RUWE$ value.

### v2

@@ 4.1 Spectra preprocessing
LAMOST is a low-resolution ($\frac{\lambda}{\Delta\lambda}\sim 1800$) spectrometer with two arms: a blue arm that covers a wavelength range of 370059003700-5900Å, and a red arm that covers a wavelength range of 570090005700-9000Å. The LAMOST survey consists of millions of spectra of stars, galaxies, and quasars. We used LAMOST DR8 and downloaded all the spectra of $A,F,G,K,M$ stars with $3000\leq T_{\mathrm{eff}}\leq 7500$K. We removed samples without measured $T_{\mathrm{eff}},\log g$ or $\mathrm{Fe}/\mathrm{H}]$ (from the LASP (Wu et al., 2014) pipeline), and samples with nonsensical negative SNR. This resulted in about 6.56.5M samples. As mentioned in Section 3.1, the hybrid approach does not require having supervised labels. Here, we decided to use only samples with labels, given the available large size of the dataset. The preprocessing of spectra is similar to what was suggested in StarGRUNet (Li & Lin, 2023): we first translate the wavelength to the rest-frame using radial velocity measurements from the LASP pipeline (Wu et al., 2014). Next, we divide the spectrum into blue and red regions and resample each region using linear interpolation. We then apply a median filter to each region, followed by a continuum normalization using a fifth-order polynomial, and a final step where we remove points higher or lower than $\pm 3\sigma$ and normalize the flux to have zero mean and unit standard deviation. For a more detailed explanation of the different pre-processing steps, please refer to Li & Lin (2023). Figure 3 shows an example of the pre-processing stages of LAMOST spectra.
@@ 4.2 Light curve preprocessing
Kepler was a space mission designed to provide high cadence light curves for stars and for the search of exoplanets through transits. As such, Kepler measured the light curves of around $200000$ main sequence and giant stars for almost $4$ years with a cadence of approx. $30$ minutes. We used Kepler samples with stellar properties from Berger et al. (2020), resulting in $183,435$ samples. The preprocessing of the light curves follows two different ways: one using the absolute magnitude and the other using the mean and standard deviation. As mentioned in 3.1.2, the reason for the two normalizations is that different properties require different light curve information. The activity and luminosity, for example, are related to the absolute amplitude of the light curve (normalized by the distance). On the other hand, the period is more easily detected using standard normalization, like zero mean and unit standard deviation. The absolute magnitude normalization was done by first dividing the light curve by $2^{-k}$ where $k$ is the absolute magnitude, calculated using the Kepler magnitude, KMAG, and the distance from Berger et al. (2020). In addition to the raw light curve, as mentioned earlier, we also calculated the FFT and the ACF, which help to determine the period. Figure 4 shows an example of the pre-processing stages of the Kepler light curve.
@@ 5 Implementation details
In big models, like DESA, hyperparameter tuning can be a very challenging task. This becomes even harder when the model consists of two steps - pre-training and alignment. We therefore chose to use a simple heuristic when defining the hyperparameters of our model. As the number of spectra samples is much larger than the number of light curve samples (6.56.5M vs 200200K), we designed the individual encoders such that the spectra encoder has more parameters than the light curve encoder. This is motivated by neural scaling laws, a phenomenological relationship between dataset size, model parameters, and performance that was originally found in the vision and language domains (Kaplan et al., 2020; Hoffmann et al., 2022), but recent works showed it also applies to astronomical data (Walmsley et al., 2024; Pan et al., 2024a). Therefore, the dimension of the final spectra features was chosen to be $2048$, and that of the light curve was chosen to be $256$. The number of parameters in the spectra and light curve encoders is about 500500M and 1111M, respectively. During hybrid training, $\lambda$ was chosen arbitrarily to be $0.5$. The embedding dimension in DualFormer was set to $256$, and the number of parameters in this module is also about 1111M. The light curve encoder was trained using a learning rate decay scheduler with the cosine annealing method. The initial learning rate is $2\cdot 10^{-5}$ decreasing to $2\cdot 10^{-6}$. All other modules were trained with a constant learning rate of $2\cdot 10^{-5}$. We trained all modules with AdamW optimizer (Loshchilov & Hutter, 2017). Lastly, we estimated the energy used to train the entire model using CodeCarbon [^1] https://codecarbon.io/ package. It is estimated to be $334$ kWh for the entire model, out of which $204$ kWh are for the pretraining stage. All the code used for training and experiments is publicly available on https://github.com/IlayMalinyak/DESA and on Zenodo (Kamai et al., 2025).
@@ 6.1 Pretraining Results
First, we present the results of the hybrid pre-training of individual modalities. As mentioned before, the spectra encoder was trained using $T_{\mathrm{eff}},\log g$ and $\mathrm{Fe}/\mathrm{H}]$ labels from the LASP pipeline. However, we test the results using labels from the high-resolution APOGEE survey (Abdurro’uf et al., 2022), at the same temperature range. This is a common practice and was done in previous works such as StarGRUNet (Li & Lin, 2023). The results on the test set of both LASP labels and APOGEE labels are shown in Figure 5. For each label, we also plot the $80\%$ and $50\%$ confidence intervals. It can be seen that the mean average error (MAE) on LASP is a factor of $\sim 2$ lower compared to APOGEE for all labels. This is reasonable given the fact that the APOGEE labels come from high-resolution spectra, while the spectra that are given to the model are low resolution. Another interesting difference between the labels is related to the prediction intervals. As mentioned in section 3.1, the prediction intervals were calibrated on a held-out validation set, different from the test set. This validation set uses the LASP labels. We see that for the LASP test set, this calibration results in the desired coverage ($80\%$ of the points are inside the $80\%$ interval, for example), while for the APOGEE labels, the coverage is lower. This is another evidence that there is a significant difference between the labels. We also compare our results with StarGRUNet. StarGRUNet (Li & Lin, 2023) used a GRU-based model to predict stellar parameters on a wide range of SNR values (SNR >5>5). While there are works with better performance for a specific range of SNR (Li et al. (2022), for example, who used $10<$ SNR <20<20), this is the state-of-the-art for a model that uses LAMOST low-resolution spectra and a wide range of SNR values.
Table 1 shows a comparison between StarGRUNet and DESA. We note that the results are very close. Specifically, our model performs slightly better than StarGRUNet for $T_{\mathrm{eff}}$ and $\mathrm{Fe}/\mathrm{H}]$, and slightly worse for $\log g$. However, it is worth mentioning that we used SNR >0>0 and StarGRUNet used SNR >5>5. This seemed to be a result of the fact that we incorporated the SNR information into the training. In Appendix A we show MAE vs different SNRs bins for $T_{\mathrm{eff}},\log g,$ and $\mathrm{Fe}/\mathrm{H}]$. We see that the sensitivity of our results to SNR is indeed better than StarGRUNet (see Figure 11 in their paper). However, there is still non-negligible sensitivity for both LASP and APOGEE labels. While it is probably impossible to remove all the dependency on the noise, this might be further investigated and improved in future papers.
@@
Model $T_{\mathrm{eff}}$ MAE (K) $\log g$ MAE (dex) $\mathrm{Fe}/\mathrm{H}]$ MAE (dex)
@@ 6.3 Binary Detection
We fine-tuned our model, with an added classification head, for a binary classification of binarity - all the binaries were assigned as class $1$, regardless of the detection method, and all the singles were assigned as class 0. As baselines, we trained the same models that were used in 6.2 as well as unimodal models. The unimodal models are simply the pretrained encoders with an added prediction head. Using the pretrained encoders is important as a sanity check - we want to see if our model is able to extract meaningful information from the combination of modalities and, specifically, better than each modality alone. In all models, we made sure that the number of trainable parameters is not less than that in our DESA model (1111M). Figure 12 compares the results using a precision-recall curve (right panel) and a True Negative Rate (TNR) vs. False Negative Rate (FNR) curve (left panel). Notice that a TNR-FNR curve can be seen as a flipped ROC curve, and therefore, the area under the curve (AUC) metric has the same meaning (with the area calculated in the opposite direction). Note that we opted for the TNR-FNR curve instead of the standard ROC curve for visual clarity. It can be seen that our model performs significantly better compared to all other models, with an AUC of $0.99$ and average precision (AP) of $1.00$. Figure 13 shows the confusion matrix of our model. It is interesting to compare it with the confusion matrices of all other models. Such a comparison revealed that, except for VicReg, all other models collapsed to predict the majority class (binary class). With the exception of the proposed DESA, VicReg was the only model able to predict some single stars correctly, but it did so with much lower precision for that class, in comparison to our model. The precision of VicReg on the single-star class is $48\%$, while our model achieved a precision of $88\%$ on that class. Confusion matrices of all other models are reported in Appendix 19.
@@ 6.4 Stellar age predicition
Model zero-shot accuracy Linear Regression $R^{2}$ BPRPBP-RP Accuracy GmagGmag Accuracy
@@
7 Conclusions
We presented DESA, a new multi-modality model for stellar astrophysics. DESA backbone consists of pre-trained unimodality encoders that were trained in a hybrid approach and shows state-of-the-art performance. The alignment module, DualFormer, is motivated by the observation that astrophysical data is unique and different compared to common multi-modality domains such as vision and NLP. We show that DESA learns a latent space that distills important physical information and can be easily transformed into meaningful diagrams. These diagrams can be used for tasks such as population analysis and outlier detection. The effectiveness of DESA is further demonstrated in various ways, including zero-shot, few-shot, and fine-tuning experiments to predict challenging and important labels such as binarity and age. DESA consistently outperforms all baselines on all experiments with (sometimes very) large margins, proving its superiority in the astronomical domain and marking a new era in data-driven astronomy where multiple measurements are efficiently combined to extract new insights. Each fine-tune task is of great physical importance and deserves an in-depth investigation and separate papers, which we plan for future work. DESA faces two primary limitations: interpretability and single-resolution modeling. Interpretability remains a challenge across machine learning models, lacking robust methods for results interpretation. Regarding resolution, DESA, while multi-modal, was trained on single surveys per modality (Kepler and LAMOST), without explicit consideration for survey-specific resolutions. The last point is related to the applicability of DESA with other surveys. In general, DESA can be applied to different pairs of surveys, such as APOGEE (Abdurro’uf et al., 2022) and TESS (Ricker et al., 2014), straightforwardly, since nothing in the suggested method itself is specific to LAMOST and Kepler. While the pre-processing is also general, there might be some practical adaptations needed in that part. For example, TESS samples are much shorter compared to Kepler. This means that it might be more challenging to crop each sample into a representative view (in that case, overlapping views can be used). Another example is that the ASPCAP pipeline in APOGEE already moves the wavelength into the rest frame, which makes this step redundant in the spectra preprocessing. However, extending DESA with multiple surveys of the same modality (e.g., APOGEE, LAMOST, TESS, and Kepler together) would require encoding the wavelength and cadence of each survey, which is one of our future directions. Addressing interpretability and multi-resolution capabilities stands as our main goal for future research. Nonetheless, DESA represents a potential groundbreaking advancement in multimodal stellar astronomy.
@@ 7 Conclusions
Table 3: Result of age prediction finetuning experiment.
@@ Appendix A Supplementary graphs
Figure 15: UMAP of all models. All points are in shaded gray, and 10 stars with pairs of spectra are colored on top. For each pair, all samples are colored with the same color and connected by a line. We can see the visual difference between DESA (upper figure) and all other models – in DESA, pairs are arranged very close to each other, implying that they have very similar embeddings. In other models, there are samples of the same star that are very distant, which implies very different embeddings.
Figure 16: Feature sensitivity study. The x-axis shows the feature index in the final features of DESA on the test set. The y-axis shows the difference between the original features and features that were generated with only spectral input (red) or only light curve input (gray). The blue circles mark indices where the lightcurve-only features are closer to the original features compared to spectra-only features.
@@
Figure 18: UMAP of DESA final embeddings. All points are in gray, and points with RUWE >1.4>1.4 are colored according to their RUWERUWE value.